OpenStudy (ttop0816):

please help! The following system of equations is _____. y = -x + 2x + 6y = 4 independent equivalent inconsistent

OpenStudy (unklerhaukus):

there is a term missing form your first equation

OpenStudy (anonymous):

y= -x + ??? there is a missing term.....

OpenStudy (ttop0816):

sorry! its -1/3x+2/3!

OpenStudy (anonymous):

ok, if you reduce your second equation, it becomes x +3y= 2, and making y SOF, it becomes y= 2/3 -x/3 which is the same as the first equation y= -1/3x +2/3. If this is the case, then the system is consistent, because we can find at least one set of solution, but not independent because if you draw the graph , the two lines are cincidental, not distict and intersecting